OpenStudy (yinshy):

What is the third term of this multiplication pattern? multiplication pattern using the formula 3 • 7n – 1 A. 1 B. 147 C. 21 D. 441

OpenStudy (photon336):

any ideas?

OpenStudy (onyxxelite):

good question let me think

OpenStudy (yinshy):

C?

OpenStudy (yinshy):

D

OpenStudy (yinshy):

its D right?

OpenStudy (photon336):

I'm really wondering if this is correct lol n = 1 I'm thinking the sequence goes like this: \[3•7(1)-1 = 20 \] \[3•7(2)-1 = 41\] @Zale101 @Zarkon

OpenStudy (yinshy):

so B?

OpenStudy (yinshy):

well can you try helping me with a different question in the mean time?

OpenStudy (photon336):

@Kainui

OpenStudy (photon336):

yeah sure

OpenStudy (yinshy):

What are the first five terms of the pattern with formula 3n + 4?

OpenStudy (yinshy):

@Photon336

OpenStudy (yinshy):

@Photon336 im putting it in here as no medals were given

OpenStudy (onyxxelite):

hello

OpenStudy (yinshy):

hey

OpenStudy (photon336):

well just insert n = 1, n= 2 and so on for instance n = 1 3(1)+4 = 12 n = 2 3(2)+4 = ?